Let \\(f(x)=x^{4}-2x^{2}\\). Find the all \\(c\\) (where \\(c\\) is the interception on the x-axis ) in the interval (-2, 2) such that \\(f\'(x)=0\\). ( Hint use Rolle\'s theorem )
1
Expert's answer
2019-09-12T09:25:40-0400

Function "f(x)=x^4-2x^2"  is continuous on the closed interval [-2, 2] and differentiable on the open interval (-2, 2).

"f(-2)=f(2)=8"

According to Rolle's theorem, there should be at least one point c in the open interval (a, b) for which f'(c)=0.

Let's find f'(x)

"f'(x)=4x^3-4x"

Therefore, we get the equation

"4x^3-4x=0"

"x(x^2-1)=0"

"x=-1;x=0;x=1"

All of these roots belong to the open interval (-2, 2), hence the answer is -1; 1; 0.